# QoS-Aware Load Balancing in the Computing Continuum via Multi-Player Bandits

### v1

@@ VII-C1 Per-Client QoS Success
A representative topology (scenario #1) is used to illustrate LB-level behavior. Fig. 5 shows the distribution of per-client QoS success percentage $\rho$. This plot shows the distribution of success ratios among all clients, indicating whether a strategy treats clients uniformly, as defined in the objective of the problem statement, or excessively favors some of them. Client QoS success with both proxy-mity configurations spans the full range from 0 to 100%, reflecting highly uneven behavior. Clients with rates close to 100% are typically single users of their closest instance, so that instance is never overloaded, while clients with rates near 0% are the ones that share the closest instance with multiple other clients, causing its overload. Due to occasional offloading to the non-nearest instance, proxy-mity 0.9 does not achieve 100% success for any client. Dec-SARSA has significantly better distribution, with most clients exceeding 85% QoS success, but 15 clients (out of 120) still fall below the target threshold of 90%, all connected to load balancers without a local instance. In contrast, QEdgeProxy fails to meet the target for only three clients, all marginally below the threshold, and exhibits an overall upward shift in the distribution, indicating more uniformly high per-client QoS satisfaction.
Figure 5: Distribution of client QoS success percentage $\rho$ (scenario #1).
@@ VII-C4 Processing Latency per Instance
Fig. 8 shows the p90 processing latency per instance in scenario #1, highlighting the impact of routing decisions on instance congestion. As expected, processing latency closely follows the request distribution in Fig. 8. Both proxy-mity variants overload instances i5 and i9, resulting in elevated latencies and QoS violations. While Dec-SARSA and QEdgeProxy achieve more balanced behavior, Dec-SARSA exhibits higher latency peaks as it occasionally overloads instances before redistributing requests. In contrast, QEdgeProxy maintains lower and more stable processing latencies by proactively distributing load across QoS-feasible instances. Processing latency must be interpreted jointly with network latency, as higher instance latency can still satisfy QoS for clients with sufficiently low network delay.
@@ VII-F Discussion and Limitations
VIII Conclusion and future work
This work addressed the challenge of ensuring per-client QoS satisfaction for latency-sensitive IoT applications in the dynamic and heterogeneous Computing Continuum (CC). We formulated decentralized load balancing in CC as a Multi-Player Multi-Armed Bandit (MP-MAB) problem with heterogeneous rewards and proposed QEdgeProxy, a lightweight distributed load balancer deployed on every CC node.
QEdgeProxy maintains a QoS pool—a dynamically updated set of service instances predicted to satisfy the given QoS requirements. The QoS pool is constructed using Kernel Density Estimation (KDE) over a sliding window of recent observations to estimate instance-level QoS success probabilities. Routing decisions combine these estimates with an adaptive $\varepsilon$-decay exploration scheme that dynamically adjusts in response to observed QoS satisfaction, and Smooth Weighted Round Robin (SWRR) routing, enabling each node to balance exploitation of high-quality instances with controlled exploration of alternatives. This design allows QEdgeProxy to avoid instance overloads, quickly identify improving or newly deployed instances, and deliver stable performance under time-varying network and workload conditions.
@@ VIII Conclusion and future work
As future work, as discussed in Section VII-F, we aim to extend our load balancer with support for multi-dimensional QoS objectives, collaborative QoS prediction, and independent service workloads, enabling more expressive QoS guarantees and end-to-end optimization across service pipelines. In addition, we plan to evaluate the approach in large-scale, heterogeneous real-world deployments. In particular, testing QEdgeProxy on mixed-device testbeds—including single-board computers such as Raspberry Pi, low-power AI accelerators, and near-edge micro-cloud nodes—would provide deeper insights into its performance under realistic hardware constraints and diverse runtime environments.

### v2

@@ VII-C1 Per-Client QoS Success
A representative topology (scenario #1) is used to illustrate LB-level behavior. Fig. 5 shows the distribution of per-client QoS success rate $\hat{\rho}$. This plot shows the distribution of success ratios among all clients, indicating whether a strategy treats clients uniformly, as defined in the objective of the problem statement, or excessively favors some of them. Client QoS success with both proxy-mity configurations spans the full range from 0 to 100%, reflecting highly uneven behavior. Clients with rates close to 100% are typically single users of their closest instance, so that instance is never overloaded, while clients with rates near 0% are the ones that share the closest instance with multiple other clients, causing its overload. Due to occasional offloading to the non-nearest instance, proxy-mity 0.9 does not achieve 100% success for any client. Dec-SARSA has significantly better distribution, with most clients exceeding 85% QoS success, but 15 clients (out of 120) still fall below the target threshold of 90%, all connected to load balancers without a local instance. In contrast, QEdgeProxy fails to meet the target for only three clients, all marginally below the threshold, and exhibits an overall upward shift in the distribution, indicating more uniformly high per-client QoS satisfaction.
Figure 5: Distribution of client QoS success rate $\hat{\rho}$ (scenario #1).
@@ VII-C4 Processing Latency per Instance
Fig. 8 shows the p90 (90th percentile) processing latency per instance in scenario #1, highlighting the impact of routing decisions on instance congestion. As expected, processing latency closely follows the request distribution in Fig. 8. Both proxy-mity variants overload instances i5 and i9, resulting in elevated latencies and QoS violations. While Dec-SARSA and QEdgeProxy achieve more balanced behavior, Dec-SARSA exhibits higher latency peaks as it occasionally overloads instances before redistributing requests. In contrast, QEdgeProxy maintains lower and more stable processing latencies by proactively distributing load across QoS-feasible instances. Processing latency must be interpreted jointly with network latency, as higher instance latency can still satisfy QoS for clients with sufficiently low network delay.
@@ VII-F Discussion and Limitations
VIII Conclusion
This work addressed the challenge of ensuring per-client QoS satisfaction for latency-sensitive IoT applications in the dynamic and heterogeneous Computing Continuum (CC). We formulated decentralized load balancing in the CC as a Multi-Player Multi-Armed Bandit (MP-MAB) problem with heterogeneous rewards and proposed QEdgeProxy, a lightweight decentralized load balancer (LB) deployed on every CC node.
QEdgeProxy maintains a QoS pool—a dynamically updated set of service instances predicted to satisfy the given QoS requirements. The QoS pool is constructed using Kernel Density Estimation (KDE) over a sliding window of recent observations to estimate instance-level QoS success probabilities. Routing decisions combine these estimates with an adaptive $\varepsilon$-decay exploration scheme that dynamically adjusts in response to observed QoS satisfaction, and Smooth Weighted Round Robin (SWRR) routing, enabling each LB to balance exploitation of high-quality instances with controlled exploration of alternatives. This design allows QEdgeProxy to avoid instance overloads, quickly identify improving or newly deployed instances, and deliver stable performance under time-varying network and workload conditions.